OpenStudy (anonymous):

Substitution method to solve systems of equations. Enter as an ordered pair. 2x-y=10 3x-2y=8

OpenStudy (anonymous):

(2x-y=10)x2 4x-2y=20 3x-2y=8 4x-3x=20-8 x=12 2(12)-y=10 y=14 (x,y)=(12,14)

OpenStudy (a_clan):

The above stated is elimination method. In substitution method, You add y on both sides of the equation (1). You get an equation for 'y'. Now, put this value of y in equation (2)